Could I please have some assistance with solving this system of linear equations in three variables using Cramer's rule?
x-2y=17
3y+z=3
-x+z=3
{(?,?,?)}

Answer by nabla(475)

1 -2 0 17
0 3 1 3
-1 0 1 3
Ok, so that is our matrix.



Now, x=
17 -2 0
3 3 1
3 0 1
_________
1 -2 0
0 3 1
-1 0 1
Let's find the denominator first since it will be the same for x, y, z.
3+2(1)=5.
Ok, now the numerator for x.
17(3)+2(3-3)+0=17*3=51.
So x=51/5.
I will let you work out y and z, but I'll give you the answer:
The solution set is {51/5, -17/5, 66/5}.
The important thing to remember with Cramer's rule is that you are replacing the column of the variable you're solving for WITH the column of constants. The denominator is always the matrix excluding the constant column. If you need any more assistance send me an E-mail at enabla@gmail.com.